# LASSI: An LLM-based Automated Self-Correcting Pipeline for Translating Parallel Scientific Codes

### v1

@@ V-A Evaluation Metrics
Tables VI and VII provide five metrics for each application code translation. The first metric, Runtime, provides the runtime of the LASSI-generated code. The second metric, $Ratio$, is defined as the runtime of the original source code in the target programming language divided by the runtime of the LASSI-generated code. If N/A is given, then the LASSI-generated code either failed to execute or its standard output did not match the expected result compared to the output of the source code.
Recognizing that code may be developed more than one way to achieve the same solution, we do not expect the code generated by LASSI to match the source code line-for-line. Nevertheless, evaluating the similarity between source and LASSI-generated codes provides valuable insights for comparing performance across LLMs. We include two string comparisons for code similarity, corresponding to metrics three and four:
@@ V-B OpenMP to CUDA Translations
Table VI lists the results for our OpenMP to CUDA translations. If an LLM could not generate code that was compiled, executed, or if the output was significantly different from the expected result, the record is marked as N/A.
We observe that $80\%$ of the translations from OpenMP to CUDA successfully generated executable code with results similar to the source HeCBench code in CUDA. This result strongly indicates the effectiveness of LASSI that incorporates a novel prompting strategy, provided domain knowledge, and automated self-correction to translate OpenMP to CUDA. Of these successful generations, we observe 78.1% execute with average runtimes within 10% or faster than the average runtime of the source CUDA code. Taking the ratio of 0.6 as a heuristic measure for reasonable similarity between codes, the pipeline generated 40.6% of the successful codes in the experimental set at this threshold or higher. Finally, the self-correction counts across all OpenMP to CUDA code translations remained quite low, with 65.6% of the trials generating executable code on the first attempt.
@@ V-C CUDA to OpenMP Translations
Table VII lists the results for our CUDA to OpenMP translations. The results strongly validate the feasibility of LASSI. Specifically, 85% of the translation samples successfully generated executable code with similar output as the source. Among these, 61.8% achieved average runtimes near or below those of the source OpenMP codes, 47.1% generated heuristically similar codes, and 55.9% generated executable code on the first attempt.
@@ V-D Discussion
We highlight two noteworthy findings from the experiments given in Tables VI and VII to shed light on the initial quality of the LASSI generated code. First, we observe a likely lower-quality generated code in Codestral’s translation of bsearch from CUDA to OpenMP, which may necessitate additional self-correcting prompts. The code similarity measures are moderate, and the translation successfully executed on the first attempt without errors requiring correction. We compared standard outputs between the original HeCBench and translated codes, confirming identical results except for reported timings. However, the average runtime of the translated code over multiple runs is 20$\times$ longer than that of the source code. Upon comparing the two codes, we noted the translated code only implements the default single thread, whereas the original source code explicitly sets 256 threads.
@@ VI Summary and Future Work
We plan to explore several extensions to LASSI for generating verifiable and more performant codes. In particular, we will integrate code verification to automatically compare expected results, with feedback incorporated into another self-correcting cycle.
@@ Acknowledgment
This material is based upon work supported by the U.S. Department of Energy, Office of Science, under contract number DE-AC02-06CH11357, at Argonne National Laboratory. Thanks to Michael Papka for help with naming the project.

### v2

@@ V-A Evaluation Metrics
Tables VI(b) and VII(b) provide five metrics for each application code translation. The first metric, Runtime, provides the runtime of the LASSI-generated code. The second metric, $Ratio$, is defined as the runtime of the original source code in the target programming language divided by the runtime of the LASSI-generated code. If N/A is given, then the LASSI-generated code either failed to execute or its standard output did not match the expected result compared to the output of the source code.
Recognizing that code may be developed in more than one way to achieve the same solution, we do not expect the code generated by LASSI to match the source code line-for-line. Nevertheless, evaluating the similarity between the source and LASSI-generated codes provides valuable insights for comparing performance across LLMs [27]. We include two string comparisons for code similarity, corresponding to metrics three and four:
@@ V-B OpenMP to CUDA Translations
Table VI(b) lists the results for our OpenMP to CUDA translations. If an LLM could not generate code that was compiled, executed, or if the output was significantly different from the expected result, the record is marked as N/A.
We observe that $80\%$ of the translations from OpenMP to CUDA successfully generated executable code with results similar to the source HeCBench code in CUDA. This result strongly indicates the effectiveness of LASSI, which incorporates a novel prompting strategy, provided domain knowledge, and automated self-correction to translate OpenMP to CUDA. Of these successful generations, we observe 78.1% execute with average runtimes within 10% or faster than the average runtime of the source CUDA code. Taking the ratio of 0.6 as a heuristic measure for reasonable similarity between codes, the pipeline generated 40.6% of the successful codes in the experimental set at this threshold or higher. Finally, the self-correction counts across all OpenMP to CUDA code translations remained quite low, with 65.6% of the trials generating executable code on the first attempt.
@@ V-C CUDA to OpenMP Translations
Table VII(b) lists the results for our CUDA to OpenMP translations. The results strongly validate the feasibility of LASSI. Specifically, 85% of the translation samples successfully generated executable code with similar output as the source. Among these, 61.8% achieved average runtimes near or below those of the source OpenMP codes, 47.1% generated heuristically similar codes, and 55.9% generated executable code on the first attempt.
@@ V-D Discussion
We highlight two noteworthy findings from the experiments given in Tables VI(b) and VII(b) to shed light on the initial quality of the code generated by LASSI. First, we observe a likely lower-quality generated code in Codestral’s translation of bsearch from CUDA to OpenMP, which may necessitate additional self-correcting prompts. The code similarity measures are moderate, and the translation successfully executed on the first attempt without errors requiring correction. We compared standard outputs between the original HeCBench and translated codes, confirming identical results except for reported timings. However, the average runtime of the translated code over multiple runs is 20$\times$ longer than that of the source code. Upon comparing the two codes, we noted the translated code only implements the default single thread, whereas the original source code explicitly sets 256 threads.
@@ VI Summary and Future Work
We plan to explore several extensions to LASSI for generating verifiable and more performant codes. In particular, we will integrate code verification to automatically compare expected results, with feedback incorporated through another self-correcting cycle.
@@ Acknowledgment
This work is supported in part by US National Science Foundation grant CCF-2413597. This material is based upon work supported by the U.S. Department of Energy, Office of Science, under contract number DE-AC02-06CH11357, at Argonne National Laboratory. Thanks to Michael Papka for helping with naming the project.